Project

General

Profile

20160503 » History » Version 13

Boyd, Steven, 04 May 2016 07:44

1 1 Boyd, Steven
h1. Ops/Comm Scheduling meeting
2
3
Location : CR11 in R3 @ 15:30 pm on Tuesday 3rd
4
5
Phone access :  ReadyTalk number 
6
US 866-740-1260
7
UK 0800-496-0576
8
Swiss 0800-70-53-51
9
Italy 800182592
10
Access Code: 2155402
11
12
h1. Agenda
13
14
* *Actions from last meeting*
15
** Generate appropriate table to store absorber pressure and material information  (Janusz) : (almost done)
16
** DS readback current requires calibration (DL) : stands until close to powerup
17
** Issues arising from last data taking run need to be identified and actioned (Boyd) : Stands
18
** Move the tracker CAENbus module higher up the chain (Craig)
19
** Get an estimate for how long it takes to wire in the chiller in (Andy)
20
21
* DC report
22 5 Boyd, Steven
* Work plan for the next week : http://micewww.pp.rl.ac.uk/attachments/6180/2016-05-03-MICE-Hall-Plan.xlsx
23 10 Boyd, Steven
* QP/QD work : http://micewww.pp.rl.ac.uk/attachments/6181/MICE_Electrical_Work_List_25APR16.docx
24 1 Boyd, Steven
* Water system work plan
25
* Bellows installation and FC cooldown
26
* Tracker hall probe installation
27
* AOB
28 4 Karadzhov, Yordan
** Prescaled trigger discussion (Online group). If you are interested, please indicate your availability in the doodle:
29 11 Karadzhov, Yordan
http://doodle.com/poll/wnkk4aevq938keb8
30 1 Boyd, Steven
31
* DONM : 10/3/2016
32 12 Boyd, Steven
h1. Minutes
33
34
* Present : Yordan, Paul H, Paul S, Mike C, Vicky B, Josef F, Coli W, Andy N, Craig M, Steve B, Pierrick H
35
36
* *Actions from last meeting*
37 13 Boyd, Steven
** Generate appropriate table to store absorber pressure and material information  (Janusz) : (almost done)
38
** DS readback current requires calibration (DL) : stands until close to powerup
39
** Issues arising from last data taking run need to be identified and actioned (Boyd) : Stands
40
** Get an estimate for how long it takes to wire in the chiller in (Andy) : Done - about 3-4 days.
41 12 Boyd, Steven
42
* DC report
43
** Focus coil won't go back in since the bellows have been switched around and now the screw hole tolerances no longer match.
44
** When John gets back from holiday the FC will be pulled out and the bellows swapped
45
** Daresbury work went well. An ongoing work plan is shown below
46
** US team came out with a program of work, some of which can be done but because the racks are not fully commissioned the
47
full program of work can't be completed.
48
** This week HPC have started to work on the compressor for the DS. They are concerned about the temperature of the air end - it's
49 1 Boyd, Steven
too hot to touch. They plan to replace this in August. Colin:Can we just buy a new one rather than replace in August? Mike: What we
50
get is effectively new but we should put an order in. Colin:we should do a comprehensive job on fixing the compressor so need to
51 13 Boyd, Steven
understand this issue. 
52
*** *Action Item*: Vicky will investigate the state of the refurbished air end to see what we get for our money.
53 12 Boyd, Steven
** Daresbury will allow the US to run into a short this week.
54
** Alan B : the 20 psi main relief valve for SSU needs replacing. Alan has a replacement but it needs installation. This can happen
55 1 Boyd, Steven
when John is back next week.
56
* Work plan for the next week : http://micewww.pp.rl.ac.uk/attachments/6180/2016-05-03-MICE-Hall-Plan.xlsx
57
** Henry: Water plan goes into period when the DSA is closed. If we turn the water back on we may be in a situation where the
58
water pipes to the conventional magnets could blow. If the DSA is closed we can't get to it. Colin: Good point. Henry: We also
59 13 Boyd, Steven
need the full ISIS running schedule spreadsheet. Colin:I've got something - I'll send it to you. 
60
** *Action Item*: Colin will put the ISIS dates for run-ups and machine work into his schedules to better reflect available access time.
61 12 Boyd, Steven
* QP/QD work 
62 13 Boyd, Steven
** Discussed during the DC report
63 1 Boyd, Steven
* Water system work plan : Colin:There is no day-by-day plan, but the plan for what is being done is understood. The plan is not going
64 13 Boyd, Steven
to (ha) plan due to delays in getting an external contractor. 
65
** *Action Item*: We will make some form of water system schedule available (Colin)
66
* Bellows installation and FC cooldown : Bellows discussed during DC report. We plan to pump the FC down ASAP and then, if happy,
67
start the cooldown.
68
* Tracker hall probe installation: Will start tomorrow. Craig and Melissa have a plan which will get the losses down to an acceptable level. This will only take a day or so.
69 12 Boyd, Steven
* AOB
70
** Prescaled trigger discussion: Yordan is organising a phone meeting to develop the prescaled trigger. For those interested there
71
is a Doodle poll for this meeting.
72
** Henry: Still looking for people to train as BLOCs. BLOC training will happen in June. An invitation will be given at the VC.
73
** Henry: Water detection in the trench? Pierrick: we have water detection in the trench but it's in the gutter and so shows water
74 13 Boyd, Steven
detection all the time. We need to come up with a proper plan. 
75
*** *Action Item*: Henry, Craig, Paul and Pierrick to look into optimising this for once and for all.
76
** Henry: Can Paul lower the target frame and check that the laser light levels haven't changed. Paul: Yes.